Factorise the expressions. 
 (lm+l)+m+1 ?????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(lm+l)+m+1
= lm+m+l+1
= m(l+1)+(l+1)
= (m+1)(l+1)

Thanks